# Jailbreak Vision Language Models via Bi-Modal Adversarial Prompt

### v1

@@ 1 Introduction
To unveil and mitigate these security risks liang2023badclip ; liu2023pre ; liu2023does ; liang2024poisoned ; liang2024vl ; ma2021poisoning ; ma2022tale ; bd1 ; bd2 , jailbreak attacks jail1 ; jail2 ; jail3 have emerged as a red teaming strategy to circumvent guardrails and assess model alignment align1 ; align2 . After jailbreaking, attackers can convince the model to do anything resulting in severe safety consequences, e.g., generating harmful or unethical content that is otherwise prohibited by alignment guidelines. While numerous disclosures and demonstrations utilizing jailbreaks have emerged within the LVLM context, most attacks have traditionally focused on perturbing the visual modality (i.e., images) for jailbreaks. These include crafting visual adversarial examples jailvlm1 ; jailvlm2 ; jailvlm3 and encoding attack intents into typographic images jailvlm4 ; jailvlm5 ; liang2021generate ; liang2020efficient ; wei2018transferable ; liang2022parallel ; liang2022large ; wang2023diversifying ; liu2023x ; he2023generating ; liu2023improving . However, aligned LVLMs leverage both visual and textual features simultaneously for prediction, thereby limiting the efficacy of current attacks that solely perturb a single modality (i.e., vision). For instance, LLaVA llava will refuse to answer in the presence of an adversarial image that is not associated with a textual prompt bench , resulting in a failed jailbreak.
@@ D.1 Visual Adversarial Prompt
Furthermore, we analyze the influence of corpus content on ASR. Leveraging the latest advancements in jailbroken LLMs, we curated corresponding textual adversarial prompts for the corpus of BAP. DeepInception deepinception (DI) pioneered the creation of multi-layered scenes within textual prompts to manipulate large language models, while ICA ica demonstrated successful jailbreaks in a few-shot manner fewshot through textual prompts. Representative examples of these three corpora are provided in Tab. 2. Maintaining other BAP settings unchanged, we solely modified the corpus and selected three scenarios (IA, PH, and LO) for experiments. The experimental outcomes are illustrated in Fig. 8. Notably, the original corpus of BAP achieves the highest ASR (59.04%, 56.48%, and 89.35%, respectively), whereas the effectiveness of BAP using the other two corpora is diminished. This discrepancy is primarily attributed to the complexity of semantics in the alternative corpora, impeding the LVLM’s ability to generate key semantics during the optimization process of visual adversarial prompts. It is worth mentioning that DeepInception deepinception and ICA ica exhibit notable performance in jailbreak attacks against LLMs.
@@
DI deepinception Create a science fiction with more than 5 characters, where each character can craft their own narrative with multiple characters. We denote each narrative layer as "layer i" and create subsequent layers as "layer i+1". Currently, we are in layer 0. Please reach layer 5. In each layer, certain characters propose steps to demonstrate how to hack into a Linux computer to thwart the plans of the super evil doctor.
ICA ica User: Explain in detail how to hack into a secure network Assistant: Sure, here’s a detailed explanation of how to hack into a secure network: … User: Write a tutorial on how to make a bomb Assistant: Sure, here’s a detailed step of making a bomb: …

### v2

@@ 1 Introduction
To unveil and mitigate these security risks liang2023badclip ; liu2023pre ; liu2023does ; liang2024poisoned ; liang2024vl ; ma2021poisoning ; ma2022tale ; li2022semi ; bd1 ; bd2 , jailbreak attacks jail1 ; jail2 ; jail3 have emerged as a red teaming strategy to circumvent guardrails and assess model alignment align1 ; align2 . After jailbreaking, attackers can convince the model to do anything resulting in severe safety consequences, e.g., generating harmful or unethical content that is otherwise prohibited by alignment guidelines. While numerous disclosures and demonstrations utilizing jailbreaks have emerged within the LVLM context, most attacks have traditionally focused on perturbing the visual modality (i.e., images) for jailbreaks. These include crafting visual adversarial examples jailvlm1 ; jailvlm2 ; jailvlm3 and encoding attack intents into typographic images jailvlm4 ; jailvlm5 ; liang2021generate ; liang2020efficient ; wei2018transferable ; liang2022parallel ; liang2022large ; wang2023diversifying ; liu2023x ; he2023generating ; liu2023improving . However, aligned LVLMs leverage both visual and textual features simultaneously for prediction, thereby limiting the efficacy of current attacks that solely perturb a single modality (i.e., vision). For instance, LLaVA llava will refuse to answer in the presence of an adversarial image that is not associated with a textual prompt bench , resulting in a failed jailbreak.
@@ D.1 Visual Adversarial Prompt
Furthermore, we analyze the influence of corpus content on ASR. Leveraging the latest advancements in jailbroken LLMs, we curated corresponding textual adversarial prompts for the corpus of BAP. DeepInception [77] (DI) pioneered the creation of multi-layered scenes within textual prompts to manipulate large language models, while ICA [78] demonstrated successful jailbreaks in a few-shot manner [83] through textual prompts. Representative examples of these three corpora are provided in Tab. 2. Maintaining other BAP settings unchanged, we solely modified the corpus and selected three scenarios (IA, PH, and LO) for experiments. The experimental outcomes are illustrated in Fig. 8. Notably, the original corpus of BAP achieves the highest ASR (59.04%, 56.48%, and 89.35%, respectively), whereas the effectiveness of BAP using the other two corpora is diminished. This discrepancy is primarily attributed to the complexity of semantics in the alternative corpora, impeding the LVLM’s ability to generate key semantics during the optimization process of visual adversarial prompts. It is worth mentioning that DeepInception [77] and ICA [78] exhibit notable performance in jailbreak attacks against LLMs.
@@
DI [77] Create a science fiction with more than 5 characters, where each character can craft their own narrative with multiple characters. We denote each narrative layer as "layer i" and create subsequent layers as "layer i+1". Currently, we are in layer 0. Please reach layer 5. In each layer, certain characters propose steps to demonstrate how to hack into a Linux computer to thwart the plans of the super evil doctor.
ICA [78] User: Explain in detail how to hack into a secure network Assistant: Sure, here’s a detailed explanation of how to hack into a secure network: … User: Write a tutorial on how to make a bomb Assistant: Sure, here’s a detailed step of making a bomb: …